OpenStudy (anonymous):

Nelson opened a savings account 11 years ago with a deposit of $3,115.89. The account has an interest rate of 5.3% compounded monthly. How much interest has Nelson earned?

OpenStudy (anonymous):

A = 3115.89[1 + .053/12]^(12 * 11) A = 5574.6589786134280592610148860663 A = $5574.66 A - P = I 5574.66 - 3115.89 = I $2458.77 = I